### v1

@@ 1 Introduction
The axioms $\mathsf{CC}$ and $\mathsf{DC}$ have a tight relationship with the notion of homotopy dimension (Definition 2.1). Informally, an “operation” in a topos is of homotopy dimension $\leq d$ if it sends $(m+d)$/̄connected maps into $m$/̄connected maps for all $m$. The axiom $\mathsf{CC}_{d}$ says that countable products are of homotopy dimension $\leq d$ (Definition 2.19). We shall see that $\mathsf{DC}_{d}$ holds in any $n$/̄topos with enough objects of homotopy dimension $\leq d$ (see Definitions 2.9 and 2.31).
@@ Acknowledgments.
We set our conventions about categories, $n$/̄topoi and connectivity. We shall say simply category to refer to an $(\infty,1)$/̄category. By default, a category is not assumed small. An $n$/̄category is a category whose mapping spaces are all $(n-1)$/̄truncated. We denote the category of (small) spaces by $\mathcal{S}$. For $-1\leq n\leq\infty$, we denote by $\mathcal{S}^{\leq n}$ the $(n+1)$/̄category of $n$/̄truncated spaces. For $n=0$, $\mathcal{S}^{\leq 0}=\mathrm{\mathcal{S}et}$ is the 1-category of sets, and for $n=\infty$, we put $\mathcal{S}^{\leq\infty}=\mathcal{S}$ by convention. An $n$/̄category $\mathcal{E}$ is an $n$/̄topos if it is an (accessible) left-exact localization of $\mathrm{Fun}\left(C,\mathcal{S}^{\leq n-1}\right)$ for a small $n$/̄category $C$ [^4] When $n<\infty$, all left-exact localizations of $\mathrm{Fun}\left(C,\mathcal{S}^{\leq n-1}\right)$ are accessible [Lur09, Remark 6.4.1.2]. This is open for $n=\infty$. An algebraic morphism (of $n$/̄topoi) $u^{*}:\mathcal{E}\to\mathcal{F}$ is a cocontinuous and left-exact functor. Such a functor always has a right adjoint $u_{*}$.
@@
For $1\leq m\leq n\leq\infty$, the $m$/̄category of $(m-1)$/̄truncated objects in an $n$/̄topos is an $m$/̄topos [Lur09, Proposition 6.4.5.4]. The $n$/̄topos envelope of an $m$/̄topos is the left adjoint to this truncation operation [Lur09, Proposition 6.4.5.7]. The original $m$/̄topos is the full subcategory of $(m-1)$/̄truncated objects of its $n$/̄topos envelope. For $\mathcal{E}$ an $\infty$/̄topos, the hypercompletion $h^{*}:\mathcal{E}\to\mathcal{E}^{\leq\infty}$ is the localization inverting the class of $\infty$/̄connected maps [Lur09, Section 6.5.2].
@@ 2.1 Homotopy dimension
    In an $n$/̄topos for $n<\infty$, every $(n-1)$/̄connected map is invertible, and so all $n$/̄topoi and algebraic morphisms between $n$/̄topoi are of homotopy dimension $\leq n$.
@@ 2.2 Countable choice
  6. f)
    The $\infty$/̄topos of sheaves on the interval $[0,1]$ does not satisfy $\mathsf{CC}_{0}$. For each $k\geq 1$, choose a cover $(U_{k,\alpha})$ of $[0,1]$ by open sets of diameter less than $1/k$, and set $A_{k}=\coprod_{\alpha}U_{k,\alpha}$. The product $\prod A_{k}$ is empty, since there is no nonempty open set $U\subseteq[0,1]$ small enough that all of the objects $A_{k}$ have sections over $U$. On the other hand, we will see below that this $\infty$/̄topos does satisfy $\mathsf{CC}_{1}$. More generally, for $0\leq d<\infty$, the $\infty$/̄topos of sheaves on the $d$/̄cube $[0,1]^{d}$ satisfies $\mathsf{CC}_{d}$ but not $\mathsf{CC}_{d-1}$; see Example 2.25.
@@
  8. h)
@@
Let us explain the intuitive meaning of the axiom $\mathsf{CC}_{d}$. The failure of $\mathsf{CC}_{0}$ can be understood as follows. Let $X$ be a topological space, $I$ a set, and $(A_{i})$ a family of inhabited sheaves on $X$ (i.e. whose stalks are all nonempty). A local section of $\prod A_{i}$ around a point $x$ is a neighborhood $U$ of $x$ and a section of every $A_{i}$ over $U$. Every $A_{i}$ has local sections around $x$, but because the neighborhoods of $x$ are only closed under finite intersections, it might not be possible to find a section of all the $A_{i}$ over the same neighborhood of $x$. For this reason, the stalk of $\prod A_{i}$ can be empty if the set $I$ is infinite, preventing the map $\prod A_{i}\to 1$ from being a surjection. Of course, this does not happen if $x$ has enough neighborhoods where surjective sheaves always have sections (i.e. if $X$ is locally of homotopy dimension $\leq 0$ around $x$, like an Alexandrov space, or the Cantor space).
The axiom $\mathsf{CC}_{1}$ says that when the $A_{i}$ are 0-connected sheaves (i.e., gerbes), then the product $\prod A_{i}$ always has local sections around $x$. However, the space of such sections might not be connected anymore. This happens if $X$ is locally of covering dimension $\leq 1$ around $x$, for example in $X=[0,1]$. The key remark is that given any neighborhood $U$ of $x$ we can always find a section of a sheaf $A$ on $U$ when $A$ is 0-connected. Essentially, one starts with local sections on a cover $U_{\alpha}$ of $U$, and because $A$ is connected, one can find homotopies between these sections on the $U_{\alpha\beta}$, and use them to construct a global section on $U$. Because $X$ is of covering dimension $\leq 1$, there are no triple intersections to consider. (If $X$ had been of covering dimension $\leq 2$, we would have needed coherence homotopies on triple intersections. Such homotopies always exist if $A$ is a 1-connected sheaf.) Given a family of connected sheaves $A_{k}$, we see that we can construct a section of each of $A_{k}$ on the same neighborhood of $x$. Thus the stalk of $\prod A_{i}$ at $x$ is not empty.
@@
Using the terminology of Remark 2.6, we can say that $\mathsf{CC}_{d}$ holds in $\mathcal{E}$ if and only if the object $\mathbb{N}\in\mathcal{E}$ is internally of homotopy dimension $\leq d$.
@@
We show that $\mathcal{E}$ does not have $\mathsf{CC}_{d-1}$. We already checked this for $d=1$ in 2.20 (f), so assume $d\geq 2$. We take $A\in\mathcal{E}$ to be the Eilenberg–Mac Lane object $K(F,d)$ for a sheaf of abelian groups $F$ on $X$ constructed below in such a way that $\pi_{0}(A^{\mathbb{N}})$ is not the terminal object of $\mathcal{E}$. Then $A$ will be $(d-1)$/̄connected, but $A^{\mathbb{N}}$ will not be 00/̄connected, so $\mathcal{E}$ does not satisfy $\mathsf{CC}_{d-1}$.
@@
Recall that the hypercompletion $h^{*}:\mathcal{E}\to\mathcal{E}^{\leq\infty}$ of an $\infty$/̄topos reflects $k$/̄connected maps for $-1\leq k\leq\infty$: a map in $\mathcal{E}$ is a $k$/̄connected map if and only if its image in $\mathcal{E}^{\leq\infty}$ is a $k$/̄connected map (see [Lur18, Proposition A.4.2.1 and Example A.4.2.4] or [ABFJ24, Proposition 4.3.2]).
@@
If $\mathsf{CC}_{d}$ holds in $\mathcal{E}$, then it also holds in its hypercompletion $\mathcal{E}^{\leq\infty}$.
@@
Let $(X_{k}\to Y_{k})$ be a countable family of $(m+d)$/̄connected maps in $\mathcal{E}^{\leq\infty}$, and let $p:X\to Y$ be their product in $\mathcal{E}^{\leq\infty}$. The map $h_{*}p:h_{*}X\to h_{*}Y$ is the product of the maps $h_{*}X_{k}\to h_{*}Y_{k}$ in $\mathcal{E}$, since $h_{*}$ preserves products. Using $h^{*}h_{*}=id$ and the reflection property of $h^{*}:\mathcal{E}\to\mathcal{E}^{\leq\infty}$, we deduce that the maps $h_{*}X_{k}\to h_{*}Y_{k}$ are $(m+d)$/̄connected maps in $\mathcal{E}$. By assumption on $\mathcal{E}$, the map $h_{*}p:h_{*}X\to h_{*}Y$ is $m$/̄connected in $\mathcal{E}$. Thus the map $h^{*}h_{*}p=p$ is $m$/̄connected in $\mathcal{E}^{\leq\infty}$. ∎
A counterexample to the converse of Proposition 2.26 will be given in Remark 4.3.

### v2

@@ 1 Introduction
The axioms $\mathsf{CC}$ and $\mathsf{DC}$ have a tight relationship with the notion of homotopy dimension (Definition 2.1). Informally, an “operation” in a topos is of homotopy dimension $\leq d$ if it sends $(m+d)$/̄connected maps into $m$/̄connected maps for all $m$. The axiom $\mathsf{CC}_{d}$ says that countable products are of homotopy dimension $\leq d$ (Definition 2.19). We shall see that $\mathsf{DC}_{d}$ holds in any $n$/̄topos with enough objects of homotopy dimension $\leq d$ (see Definitions 2.9 and 2.37).
@@ Acknowledgments.
We set our conventions about categories, $n$/̄topoi and connectivity. We shall say simply category to refer to an $(\infty,1)$/̄category. By default, a category is not assumed small. An $n$/̄category is a category whose mapping spaces are all $(n-1)$/̄truncated. We denote the category of (small) spaces by $\mathcal{S}$. For $-1\leq n\leq\infty$, we denote by $\mathcal{S}^{\leq n}$ the $(n+1)$/̄category of $n$/̄truncated spaces. For $n=0$, $\mathcal{S}^{\leq 0}=\mathrm{\mathcal{S}et}$ is the 1-category of sets, and for $n=\infty$, we put $\mathcal{S}^{\leq\infty}=\mathcal{S}$ by convention. An $n$/̄category $\mathcal{E}$ is an $n$/̄topos if it is an (accessible) left-exact localization of $\mathrm{Fun}\left(C,\mathcal{S}^{\leq n-1}\right)$ for a small $n$/̄category $C$. (When $n<\infty$, all left-exact localizations of $\mathrm{Fun}\left(C,\mathcal{S}^{\leq n-1}\right)$ are accessible [Lur09, Remark 6.4.1.2]. This is open for $n=\infty$.) An algebraic morphism (of $n$/̄topoi) $u^{*}:\mathcal{E}\to\mathcal{F}$ is a cocontinuous and left-exact functor. Such a functor always has a right adjoint $u_{*}$.
@@
For $1\leq m\leq n\leq\infty$, the $m$/̄category of $(m-1)$/̄truncated objects in an $n$/̄topos is an $m$/̄topos [Lur09, Proposition 6.4.5.4]. The $n$/̄topos envelope of an $m$/̄topos is the left adjoint to this truncation operation [Lur09, Proposition 6.4.5.7]. The original $m$/̄topos is the full subcategory of $(m-1)$/̄truncated objects of its $n$/̄topos envelope. For $\mathcal{E}$ an $\infty$/̄topos, the hypercompletion $P_{\infty}:\mathcal{E}\to\mathcal{E}^{\leq\infty}$ is the localization inverting the class of $\infty$/̄connected maps [Lur09, Section 6.5.2].
@@ 2.1 Homotopy dimension
    In an $n$/̄topos for $n<\infty$, every $(n-1)$/̄connected map is invertible, and so all $n$/̄topoi and algebraic morphisms between $n$/̄topoi are of homotopy dimension $\leq n$. (Beware the $m$-topos envelope of an $n$-topos, for $m>n$, may have a strictly larger homotopy dimension.)
@@ 2.2 Countable choice
    By Lemma 2.8 (4), if $\mathsf{CC}_{d}$ holds in $\mathcal{E}$, then it holds in every slice of $\mathcal{E}$.
  3. c)
@@
    The $\infty$/̄topos of sheaves on the interval $[0,1]$ does not satisfy $\mathsf{CC}_{0}$. For each $k\geq 1$, choose a cover $(U_{k,\alpha})$ of $[0,1]$ by open sets of diameter less than $1/k$, and set $A_{k}=\coprod_{\alpha}U_{k,\alpha}$. The product $\prod A_{k}$ is empty, since there is no nonempty open set $U\subseteq[0,1]$ small enough such that all of the objects $A_{k}$ have sections over $U$. On the other hand, we will see below that this $\infty$/̄topos does satisfy $\mathsf{CC}_{1}$. More generally, for $0\leq d<\infty$, the $\infty$/̄topos of sheaves on the $d$/̄cube $[0,1]^{d}$ satisfies $\mathsf{CC}_{d}$ but not $\mathsf{CC}_{d-1}$; see Example 2.26.
  8. h)
@@
  9. i)
    We will see in Proposition 2.27 that if an $\infty$/̄topos $\mathcal{E}$ has $\mathsf{CC}_{d}$, then so does the $n$-topos $\mathcal{E}^{\leq n-1}$ for any $n$.
  10. j)
@@
Let us explain intutively the geometric meaning of the axiom $\mathsf{CC}_{d}$. The failure of $\mathsf{CC}_{0}$ can be understood as follows. Let $X$ be a topological space, $I$ a set, and $(A_{i})$ a family of inhabited sheaves on $X$ (i.e. whose stalks are all nonempty). A local section of $\prod A_{i}$ around a point $x$ is a neighborhood $U$ of $x$ and a section of every $A_{i}$ over $U$. Every $A_{i}$ has local sections around $x$, but because the neighborhoods of $x$ are only closed under finite intersections, it might not be possible to find a section of all the $A_{i}$ over the same neighborhood of $x$. For this reason, the stalk of $\prod A_{i}$ can be empty if the set $I$ is infinite, preventing the map $\prod A_{i}\to 1$ from being a surjection. Of course, this does not happen if $x$ has enough neighborhoods where surjective sheaves always have sections (i.e. if $X$ is locally of homotopy dimension $\leq 0$ around $x$, like an Alexandrov space, or the Cantor space).
The axiom $\mathsf{CC}_{1}$ says that when the $A_{i}$ are 0-connected sheaves (i.e., gerbes), then the product $\prod A_{i}$ always has local sections around $x$. However, the space of such sections might not be connected anymore. This happens if $X$ is locally of covering dimension $\leq 1$ around $x$, for example in $X=[0,1]$. The key remark is that given any neighborhood $U$ of $x$, we can always find a section of a sheaf $A$ on $U$ when $A$ is 0-connected. Essentially, one starts with local sections on a cover $U_{\alpha}$ of $U$, and because $A$ is connected, one can find homotopies between these sections on the $U_{\alpha\beta}$, and use them to construct a global section on $U$. Because $X$ is of covering dimension $\leq 1$, there are no triple intersections to consider. (If $X$ had been of covering dimension $\leq 2$, we would have needed coherence homotopies on triple intersections. Such homotopies always exist if $A$ is a 1-connected sheaf.) Given a family of connected sheaves $A_{k}$, we see that we can construct a section of each of $A_{k}$ on the same neighborhood of $x$. Thus the stalk of $\prod A_{i}$ at $x$ is not empty.
@@
We now discuss the logical meaning of the axiom $\mathsf{CC}_{d}$. Recall first that, in ordinary logic, the class of surjections is used to provide semantics for existential conditions. Using the notation of [HoTT], a proposition $\exists x.P(x)$ is interpreted as the image of the map $p:\Sigma_{x:X}P(x)\to 1$, that is as the maximal subterminal object over which the map $p$ is a surjection. The class of isomorphisms is similarly used for unique existential conditions. The proposition $\exists!x.P(x)$ is interpreted as the condition that the map $p:\Sigma_{x:X}P(x)\to 1$ is an isomorphism, i.e., the maximal subterminal object over which the map $p$ is an isomorphism.
Between these two conditions, the classes of $n$-connected maps provide semantics for an (external) hierarchy of intermediate “uniqueness levels” for existential conditions. Let us introduce the notation $\exists^{n}x.P(x)$ to mean that $\Sigma_{x:X}P(x)$ is an $n$-connected object. Informally, for $n\geq 0$, $\exists^{n}x.P(x)$ means that there exists $x$ satisfying $P(x)$ and for any two such $x$, there exists a path (or equality) between them in the sense of $\exists^{n-1}$. Semantically, $\exists^{n}x.P(x)$ corresponds to the maximal subterminal object over which the map $p$ is $n$-connected. When $n=-1$, this recovers the quantifier $\exists$. When $n=\infty$, this does not recover the quantifier $\exists$!∃ !, since the semantics of $\exists^{\infty}x.P(x)$ is that the object $\Sigma_{x:X}P(x)$ is $\infty$/̄connected rather than contractible.
$\exists!\ \Rightarrow\ \exists^{\infty}\ \Rightarrow\ \dots\ \Rightarrow\ \exists^{0}\ \Rightarrow\ \exists^{-1}\,=\,\exists\,.$
With these notions in mind, any operation sending $(m+d)$-connected maps to $m$-connected maps (like $\prod_{\mathbb{N}}$ in $\mathsf{CC}_{d}$) can be thought of as an operation “weakening the uniqueness level” of existential conditions.
Remark 2.23.
Using the terminology of Remark 2.6, we can say that $\mathsf{CC}_{d}$ holds in $\mathcal{E}$ if and only if the object $\mathbb{N}\in\mathcal{E}$ is internally of homotopy dimension $\leq d$. This condition can also be expressed in the language of homotopy type theory, for instance as
$\mathsf{CC}^{\mathsf{int}}:\mathbb{N}\to\mathsf{Prop}\,,\quad\qquad\mathsf{CC}_{d}^{\mathsf{int}}:=\prod_{X:\mathbb{N}\to\mathcal{U}}\big{(}\prod_{n:\mathbb{N}}\mathsf{isConn}_{d-1}(X(n))\big{)}\to\mathsf{isConn}_{-1}\big{(}\prod_{n:\mathbb{N}}X(n)\big{)}\,.$
Here $\mathsf{isConn}_{m}(A)$ is the proposition that a type $A$ is $m$-connected, i.e., that the $m$-truncation $\lVert A\rVert_{m}$ is contractible. We set $\mathsf{isConn}_{\infty}(A):=\prod_{m:\mathbb{N}}\mathsf{isConn}_{m}(A)$. The statement $\mathsf{CC}_{d}^{\mathsf{int}}$ is the appropriate formulation of $\mathsf{CC}_{d}$ in an elementary $\infty$/̄topos [nLa21, Ras18]. (In this context, the external statement of Definition 2.19 may not even make sense since elementary topoi may not have external countable products when their natural number object is not standard. For the same reason, the internal definition of $\mathsf{isConn}_{\infty}$ may be stronger than its external counterpart.)
@@
We show that $\mathcal{E}$ does not have $\mathsf{CC}_{d-1}$. We already checked this for $d=1$ in 2.20 (g), so assume $d\geq 2$. We take $A\in\mathcal{E}$ to be the Eilenberg–Mac Lane object $K(F,d)$ for a sheaf of abelian groups $F$ on $X$ constructed below in such a way that $\pi_{0}(A^{\mathbb{N}})$ is not the terminal object of $\mathcal{E}$. Then $A$ will be $(d-1)$/̄connected, but $A^{\mathbb{N}}$ will not be 00/̄connected, so $\mathcal{E}$ does not satisfy $\mathsf{CC}_{d-1}$.
@@
If $\mathsf{CC}_{d}$ holds in $\mathcal{E}$, then it also holds in every $(n+1)$-topos $\mathcal{E}^{\leq n}$ for $0\leq n\leq\infty$ (in particular, in the hypercompletion $\mathcal{E}^{\leq\infty}$).
The proof of Proposition 2.27 will need a few lemmas about connected maps. In this proof, we write $P_{k}=(-)^{\leq k}:\mathcal{E}\to\mathcal{E}^{\leq k}$ for the left adjoint to the inclusion $\mathcal{E}^{\leq k}\subseteq\mathcal{E}$ for any $0\leq k\leq\infty$; $P_{\infty}:\mathcal{E}\to\mathcal{E}^{\leq\infty}$ is the hypercompletion functor.
Lemma 2.28.
If $f:X\to Y$ and $g:Y\to Z$ are two maps such that $g$ is $(k+1)$-connected and $gf$ is $k$-connected, then $f$ is $k$-connected.
@@
We consider the pullback of $g$ along $gf$
[…]
The map $Y^{\prime}\to Y$ is $k$-connected since it is a base change of $gf$. The map $Y^{\prime}\to X$ is $(k+1)$-connected since it is a base change of $g$. The map $X\to Y^{\prime}$ is $(k+1)$-connected since it is a section of a $k$-connected map (see [Lur09, Proposition 6.5.1.20] or [HoTT, Lemma 7.5.11]). The map $f$ is the composition of two $k$-connected maps, thus $k$-connected. ∎
Lemma 2.29.
If a map $X\to Y$ is sent to an isomorphism by $P_{k+1}:\mathcal{E}\to\mathcal{E}^{\leq k+1}$, then it is $k$-connected.
Proof.
Apply Lemma 2.28 to the commutative triangle $X\to Y\to P_{k+1}X=P_{k+1}Y$. ∎
Lemma 2.30.
For $k+1\leq n\leq\infty$, a map $X\to Y$ in $\mathcal{E}$ is $k$-connected if and only if the map $P_{n}X\to P_{n}Y$ is $k$-connected in $\mathcal{E}^{\leq n}$. (In other words, the functor $P_{n}:\mathcal{E}\to\mathcal{E}^{\leq n}$ preserves and reflects $k$/̄connected maps. Note that we may have $k=n=\infty$.)
Proof.
The factorization system ($k$-connected, $k$-truncated) is defined both in $\mathcal{E}$ and in $\mathcal{E}^{\leq n}$. The inclusion $\mathcal{E}^{\leq n}\subseteq\mathcal{E}$ preserves diagonals and isomorphisms, thus $k$-truncated maps. (Since it is moreover a conservative functor, it reflects $k$-truncated maps; we will use this below.) Consequently, its left adjoint $P_{n}:\mathcal{E}\to\mathcal{E}^{\leq n}$ preserves $k$-connected maps for every $k$.
Conversely, we assume that the map $P_{n}X\to P_{n}Y$ is $k$-connected in $\mathcal{E}^{\leq n}$. Let us show that it is also $k$-connected when viewed as a map in $\mathcal{E}$. We consider its ($k$-connected, $k$-truncated) factorization $P_{n}X\to Z\to P_{n}Y$ in $\mathcal{E}$. The map $Z\to P_{n}Y$ is $k$-truncated in $\mathcal{E}$, thus $n$-truncated since $k<n$. This shows that $Z$ is an object of $\mathcal{E}^{\leq n}$. Since the inclusion $\mathcal{E}^{\leq n}\subseteq\mathcal{E}$ reflects $k$-truncated maps, the map $Z\to P_{n}Y$ is also $k$-truncated in $\mathcal{E}^{\leq n}$. Since $Z$ is in $\mathcal{E}^{\leq n}$, the image of $P_{n}X\to Z$ by $P_{n}$ is itself. Since it is $k$-connected in $\mathcal{E}$, it is $k$-connected in $\mathcal{E}^{\leq n}$ by the argument above. Since $P_{n}X\to Z$ and $P_{n}X\to P_{n}Y$ are $k$-connected in $\mathcal{E}^{\leq n}$, so is $Z\to P_{n}Y$ by cancellation. This shows that $Z\to P_{n}Y$ is both $k$-connected and $k$-truncated in $\mathcal{E}^{\leq n}$, thus an isomorphism. Hence we have shown the map $P_{n}X\to P_{n}Y$ is $k$-connected in $\mathcal{E}$.
We now consider the canonical square in $\mathcal{E}$
${X}$${P_{n}X}$${Y}$${P_{n}Y\,.}$
The maps $Y\to P_{n}Y$ is $n$-connected, thus $(k+1)$-connected. The maps $X\to P_{n}X$ is $n$-connected, thus $k$-connected. The map $P_{n}X\to P_{n}Y$ is $k$-connected, so the composition $X\to P_{n}X\to P_{n}X$ is $k$-connected. Then the map $X\to Y$ is $k$-connected by Lemma 2.28. ∎
Remark 2.31.
For $n<\infty$, since not every map between $n$-connected objects is an $n$-connected map, the functor $P_{n}:\mathcal{E}\to\mathcal{E}^{\leq n}$ inverts some maps which are not $n$-connected. Since the $n$-connected maps in $\mathcal{E}^{\leq n}$ are exactly the isomorphisms, this shows that Lemma 2.30 is false when $k+1>n$.
Remark 2.32.
For $n=0$, Lemma 2.30 recovers that $X\to Y$ is a surjection if and only if the map $P_{0}X\to P_{0}Y$ is a surjection. For $n=\infty$, the result is [Lur18, Proposition A.4.2.1 and Example A.4.2.4] or [ABFJ24, Proposition 4.3.2].
Proof of Proposition 2.27.
If $n<\infty$, we have seen that $\mathsf{CC}_{d}$ holds in the $(n+1)$-topos $\mathcal{E}^{\leq n}$ for $d\geq n+1$. So, we may assume that $d\leq n$.
Let $(X_{k}\to Y_{k})$ be a countable family of $(d-1)$/̄connected maps in $\mathcal{E}^{\leq n}$, and let $p:X\to Y$ be their product in $\mathcal{E}^{\leq n}$. We denote by $\iota_{n}$ the inclusion $\mathcal{E}^{\leq n}\subseteq\mathcal{E}$. The map $\iota_{n}p:\iota_{n}X\to\iota_{n}Y$ is the product of the maps $\iota_{n}X_{k}\to\iota_{n}Y_{k}$ in $\mathcal{E}$, since $\iota_{n}$ preserves products. Using $P_{n}\iota_{n}=id$ and the reflection of $(d-1)$-connected maps of $P_{n}:\mathcal{E}\to\mathcal{E}^{\leq n}$ (Lemma 2.30), we deduce that each map $\iota_{n}X_{k}\to\iota_{n}Y_{k}$ is $(d-1)$/̄connected in $\mathcal{E}$. By assumption on $\mathcal{E}$, the map $\iota_{n}p:\iota_{n}X\to\iota_{n}Y$ is surjective in $\mathcal{E}$. Thus the map $P_{n}\iota_{n}p=p$ is surjective in $\mathcal{E}^{\leq n}$. ∎
A counterexample to the converse of Proposition 2.27 for $n=\infty$ will be given in Remark 4.3.